OpenStudy (anonymous):

how to solve this inequalities x2-7x-8>0

OpenStudy (john_es):

First, you should try to factorize the second grade polinomy, \[x^2-7x-8=(x+1)(x-8)\] Then you can solve this, \[(x+1)(x-8)>0\] looking for the sign of this equation in this way, \[\text{if}\ \ x<-1\Rightarrow -\cdot - =+\] \[\text{if}\ \ 8>x>-1\Rightarrow +\cdot - =-\] \[\text{if}\ \ x>8\Rightarrow +\cdot + =+\] So the only region where the equation is <0 is the interval (-1,8).